# SuperRL: Reinforcement Learning with Supervision to Boost Language Model Reasoning

### v1

@@ Appendix A Limitations and Future Work
Despite the strong empirical results of SuperRL, several limitations remain that point toward fruitful directions for future research.
Actor–Critic Integration Scope.
Our current SuperRL framework fuses supervised and reinforcement signals solely at the actor level, leaving the critic architecture unchanged. This limits the potential benefits of supervision on value estimation, especially in reward-sparse settings where critic instability is a known issue. While the actor benefits from adaptive loss weighting, the critic may still suffer from high variance or misaligned value targets. Future work could explore critic-aware hybridization—such as incorporating auxiliary imitation-based value heads, distillation from teacher critics trained on demonstrations, or uncertainty-guided target shaping—to jointly stabilize actor and critic learning.
Static Fusion Strategy.
The uncertainty-weighted hybrid loss in the Hybrid Actor in SuperRL provides a principled and adaptive mechanism for balancing objectives, but remains fixed across datasets once initialized. In practice, certain datasets (e.g., highly structured or low-noise) may admit more lightweight or specialized fusion strategies that require less computational overhead or training complexity. Designing dataset-aware or task-specific hybridization mechanisms—such as confidence-triggered SFT modulation or curriculum-guided loss annealing—could yield more efficient and tailored learning strategies.
Rule-Based Framework Design.
The SuperRL currently employs a hard-coded, rule-based mechanism to estimate dataset difficulty and select training strategies in the Adaptive Switch. While effective in our experiments, this approach is coarse-grained and may fail to capture nuanced distinctions in task complexity or reward structure. In future work, we plan to replace this module with a learned policy selector that dynamically predicts the optimal training configuration based on early-stage learning signals (e.g., reward variance, KL divergence trends, early plateau detection). This direction may benefit from techniques in meta-learning, reinforcement curriculum learning, or Bayesian optimization.
Scalability and Efficiency.
Although we demonstrate the broad applicability of SuperRL across multiple actor–critic algorithms and base models, its reliance on dual-loss computation and uncertainty learning introduces overhead in both compute and memory. Scaling the framework to very large models or long-horizon tasks may therefore require architectural optimizations such as parameter sharing across objectives, loss decoupling, or batch reweighting heuristics. Additionally, efficient parallelization strategies for joint SFT and RL processing are important for deployment at production scale.
In summary, while SuperRL provides a robust foundation for integrating supervised priors into reinforcement learning for language models, we believe further refinement of actor–critic design, adaptivity mechanisms, and system scalability will be crucial for advancing the field toward more generalizable and efficient hybrid training strategies.
@@ Appendix B Broader Impacts and Safeguards
The proposed SuperRL framework introduces a new hybrid reinforcement learning strategy that integrates SFT loss into the actor’s update loop. This approach may carry several broader impacts, particularly in the context of increasingly capable and general-purpose language models. On the one hand, the improvements in convergence stability, sample efficiency, and generalization under sparse-reward conditions make this method attractive for domains where reinforcement learning has traditionally struggled. These include education, programming assistance, automated scientific discovery, and other settings where precise reasoning and data efficiency are critical. The reduced reliance on large volumes of reward-labeled data may also lower the entry barrier for fine-tuning smaller models in more constrained or specialized environments.
At the same time, as with many advances in RL and LLM training, potential concerns arise regarding misuse and unintended consequences. Enhanced reasoning capabilities may inadvertently be applied in domains such as persuasive content generation, misinformation synthesis, or manipulation of human preferences—especially if downstream applications fail to incorporate adequate human oversight. Furthermore, the incorporation of supervised signals throughout training means that any biases embedded in the training corpus may persist or even be reinforced during policy updates, particularly if the model overfits to surface-level heuristics present in expert demonstrations.
It is also worth noting that while the hybrid loss improves stability, it does not guarantee robustness or fairness under distributional shifts. The method assumes access to high-quality expert data and reasonably aligned reward signals; in settings where these assumptions do not hold, model behavior may become unpredictable or brittle. In addition, the hybrid approach may make model attribution more difficult, as it intertwines supervised and reinforcement-driven behaviors during learning, complicating post hoc interpretation of model decisions.
We do not make specific claims regarding the societal or safety alignment of models trained under this framework. The practical implications of SuperRL —positive or negative—will ultimately depend on the nature of its application and the surrounding deployment context. Nonetheless, we believe that discussions around the responsible development and use of hybrid training frameworks are important and ongoing. Future research may benefit from more targeted evaluations of fairness, robustness, and transparency in hybrid optimization, especially as such techniques become more widely adopted in real-world systems.
@@ C.1 Dataset Configurations
To rigorously evaluate SuperRL, we utilize six diverse datasets encompassing various reasoning challenges, from arithmetic problem-solving to complex symbolic logic and semi-structured data analysis. Below, we provide detailed descriptions of each dataset:
@@ C.5 Reward Design
  • Performing symbolic equivalence checks (e.g., $42x+$ 4 vs. $4+2x$);
@@ C.7 Environment Setup
For rollout generation, we employ the vLLM backend, which enables efficient and scalable batched decoding with optimized GPU memory usage. Each prompt is decoded to produce five candidate responses, allowing for diverse sampling during training. To ensure stable runtime behavior, we cap GPU memory utilization at 40%. The decoding configuration adopts a stochastic sampling strategy with a temperature of 1.0, top-$k=-1$ (disabled), and top-$p=1.0$, corresponding to unconstrained nucleus sampling. These settings encourage diverse yet coherent response generation from the model. We constrain the maximum sequence length—comprising both the input prompt and the generated output—to 2048 tokens. Prompts that exceed this limit are filtered out prior to generation, and any attempt to exceed the limit during decoding raises a truncation error. This strict enforcement helps maintain consistency and prevents the introduction of ambiguous or malformed training signals.
@@
C.8 Details for K-step Parameter Settings Exploration
We conduct ablation experiments using four initialization settings for the uncertainty weights: $\sigma_{pg}=1,\sigma_{sft}=1$, $\sigma_{pg}=1,\sigma_{sft}=0$, $\sigma_{pg}=0,\sigma_{sft}=1$, and our default $\sigma_{pg}=0,\sigma_{sft}=0$, denoted respectively as settings 11, 10, 01, and 00. We apply Hybrid training on GSM8k using QWen2.5-1.5B, and the Figure 4 5 show the evolution of scores on GSM8k and PRM12k as training steps increase.
Figure 4: Validation Score on GSM8k When Trained on GSM8k Using Different Initial Strategies
Figure 5: Val Score on PRM12k When Trained on GSM8k Using Different Initial Strategies
We observe that on GSM8k, all four settings lead to similar performance trajectories, suggesting that this task is relatively insensitive to the relative weighting between policy gradient and imitation loss. The reward signal alone is strong enough to guide effective learning. In contrast, on PRM12k, the model collapses under the 10 setting ($\sigma_{pg}=1,\sigma_{sft}=0$), likely because the absence of supervision during initial training harms the model’s ability to generalize to more difficult PRM12k—whose distribution differs significantly and exhibits stronger structural regularities. The 00 setting ($\sigma_{pg}=0,\sigma_{sft}=0$) achieves the best performance across both tasks, indicating that assigning strong and balanced weights to both the policy gradient and imitation losses at the beginning of training contributes to effective exploration and generalization.
[…]
The following table shows representative examples over the first 10 RL steps using QWen2.5-1.5B trained with GRPO on GSM8K and HiTab. Each entry reports the average reward at a specific step, along with the total number of increases between adjacent steps.
Dataset Step1 Step2 Step3 Step4 Step5 Step6 Step7 Step8 Step9 Step10 $increase\_num$ $recent\_avg\_reward$
GSM8K 0.0617 0.0570 0.0828 0.1305 0.1820 0.2398 0.3148 0.3695 0.4609 0.4984 10 0.2397
HiTab 0.0008 0.0008 0.0023 0 0 0 0 0 0 0 1 0.0031
Table 7: Step-wise average reward during the first 10 RL updates on GSM8K and HiTab using QWen2.5-1.5B ($batch\_size=256$). Steps are split into two blocks for compact readability.
@@ Appendix D Lessons Learned from Failed Attempts and Design Iterations
D.1 The Pitfalls of Integrating Online Rollouts into Supervised Fine-Tuning in Hybrid Actor
We explored an on-policy training strategy that integrates supervised learning directly Hybrid Actor by extracting learning signals from rollout responses(for GRPO algorithm). Unlike naive self-imitation, our method uses a reward-aware filtering mechanism: rollout responses are adopted as supervision targets only if they receive non-zero rewards; otherwise, the model reverts to offline expert-annotated responses. This allows the training signal to be dynamically chosen based on the quality of exploration outcomes.This mechanism enables the actor to optimize toward both policy improvement and imitation from valid behavior traces.
Despite the theoretical appeal of this formulation, its performance on reward-sparse benchmarks was consistently poor. As shown in Table 8, all variants of this reward-aware rollout-to-SFT approach achieved near-zero test scores on LIMO, a dataset characterized by complex multi-step reasoning and weak reward density.
Experimental Results. Despite its theoretical appeal, this rollout-to-SFT strategy failed to yield robust generalization in practice. As summarized in Table 8, all evaluated runs on the LIMO benchmark—representing complex, long-horizon reasoning tasks—collapsed to near-zero accuracy. This result holds across variations in random seed, training duration, and log-sigma scheduling.
Table 8: On-Policy Evaluation Results on LIMO (Rollout-to-SFT)
Run Name Final Score Min Score Max Score
Qwen2.5-1.5B_20250502 0.000 0.000 0.043
Qwen2.5-1.5B_20250501 0.000 0.000 0.042
Qwen2.5-1.5B_20250430 0.000 0.000 0.018
Generalization Across Datasets. This failure pattern is not unique to the LIMO dataset. Across other reasoning benchmarks such as PRM12K, OpenR1, and HiTab, we observed similarly unstable training dynamics under the rollout-to-SFT paradigm. In some cases, such as PRM12K, the model exhibits a short-lived improvement in validation accuracy during early training, only to collapse irreversibly in later stages—indicative of an overfitting loop to spurious patterns in reward-positive trajectories. In others, including OpenR1 and HiTab, the model degrades steadily throughout training, failing to extract any durable learning signal from the reward-conditioned responses. Regardless of the specific trajectory, all runs eventually converge to near-zero accuracy, further affirming the brittleness of this approach in sparse or deceptive reward landscapes.
Failure Analysis. We identify several factors contributing to this consistent collapse. First, the strategy relies on the current policy to generate both inputs and supervision targets, which introduces a strong confirmation bias—especially problematic when rewards are sparse or noisy. The model reinforces its own suboptimal behavior without sufficient external correction, leading to overfitting on early or lucky successes. Second, since reward filtering allows only high-reward responses to propagate gradients, the effective batch size of learning signals becomes highly unstable and often too small for reliable policy updates. Third, the fallback to offline supervision in zero-reward cases introduces distributional mismatch between positive and negative samples, creating an inconsistent optimization target that destabilizes convergence. Finally, by tightly coupling exploration and imitation in an on-policy loop, the model becomes more susceptible to cascading errors—where minor deviations early in training can quickly spiral into mode collapse under the compounding effects of self-reinforcement.
Collectively, these findings suggest that while reward-aware rollout-to-SFT is conceptually appealing for integrating exploration and learning, its naive implementation fails to provide a stable or scalable training signal in practice—particularly under the reward sparsity and multi-step dependencies typical of reasoning-centric tasks.
D.2 Offline Data as GRPO-Compatible Rollouts: A Negative Result
@@
D.2.1 Variant I: Direct Injection of Expert Rollouts
@@
D.2.2 Variant II: Self-Rewritten Expert Rollouts
@@
D.3 Conclusion: Offline Rollouts Undermine GRPO Stability
@@
D.4 Offline Data as Few-shots

### v2

@@ Appendix A Limitations and Future Work
While SuperRL demonstrates strong empirical performance across a variety of reasoning benchmarks, several limitations remain. First, the current fallback mechanism is reactive—it only resorts to supervised fine-tuning (SFT) when all sampled trajectories receive zero reward. While this avoids overwriting useful gradient signals from positive rollouts, it does not proactively leverage the offline data to shape early-stage exploration or mitigate suboptimal trajectory distributions. Incorporating more proactive or uncertainty-aware fallback strategies—such as reward prediction confidence, entropy thresholds, or offline pre-filtering—could further enhance learning stability.
Second, although SuperRL dynamically interleaves RL and SFT at the instance level, the two optimization objectives remain loosely coupled, alternating between distinct update rules. This decoupling may result in suboptimal credit assignment and hinder smoother gradient integration. A promising direction is to develop fully unified objectives or joint loss formulations that softly interpolate between SFT and RL signals, possibly guided by reward magnitude or rollout quality.
Third, our framework currently assumes access to high-quality expert trajectories for fallback SFT. However, such data may be unavailable or limited in some domains. Future work may explore bootstrapping fallback supervision from model self-refinements, synthetic generation, or retrieval-augmented reasoning.
Lastly, SuperRL is evaluated on static reasoning tasks with fixed inputs. Extending the framework to interactive or agentic environments—where reasoning is coupled with tool use, memory, or long-horizon planning—would require addressing new challenges in credit assignment, non-stationarity, and temporal abstraction. Moreover, integrating SuperRL with test-time search or planning methods (e.g., MCTS, beam-guided rollouts) remains an exciting avenue to push the limits of adaptive reasoning under sparse rewards.
Overall, while SuperRL provides a practical and effective solution to unify supervised and reinforcement-based reasoning, we believe it opens a broader design space for adaptive training paradigms that blend exploration, supervision, and search.
@@ Appendix B Broader Impacts and Safeguards
As language models become increasingly capable of performing complex reasoning tasks, they are poised to influence high-stakes domains such as education, scientific discovery, legal analysis, and decision support. SuperRL, by improving the reasoning robustness of LLMs, could further accelerate this trend. However, such capabilities also introduce new societal risks, especially if models generate persuasive yet flawed reasoning or learn undesirable behaviors from biased feedback.
A key benefit of SuperRL lies in its fallback to high-quality supervised trajectories when reward signals are absent. This acts as a built-in safeguard, anchoring learning to human-verified reasoning paths and reducing the likelihood of reinforcement-induced divergence. However, reliance on offline data also inherits its limitations—such as annotation bias, narrow coverage, or stylistic homogenization—which could be amplified by repeated reuse. Careful curation, diversity analysis, and auditing of demonstration datasets are therefore crucial to mitigate these effects.
Moreover, since SuperRL adaptively integrates reinforcement learning, it inherits known risks from RL-based training: reward hacking, undesired generalization, and non-transparent credit assignment. Although our instance-level switching reduces over-optimization on spurious reward signals, stronger safeguards—such as adversarial evaluation, uncertainty estimation, or reward model interpretability—are needed to ensure alignment in more open-ended settings.
From a deployment perspective, reasoning models trained with SuperRL should be clearly scoped and audited before use in sensitive applications. Developers should monitor not only final accuracy but also behavioral changes introduced by the RL component, especially in failure cases. Human-in-the-loop evaluation, scenario red-teaming, and counterfactual probing can provide additional oversight.
Lastly, while SuperRL improves reasoning competence, it does not address deeper epistemic limitations of LLMs, such as lack of verifiability or causal grounding. Future safeguards may benefit from integrating structured verification, external tool use, or modular reasoning pipelines to augment the transparency and controllability of these systems.
@@ C.1 Dataset Configurations
To rigorously evaluate SuperRL, we utilize nine diverse datasets encompassing various reasoning challenges, from arithmetic problem-solving to competition-level mathematics and multi-agent reasoning. Below, we provide detailed descriptions of each dataset:
@@
AIME 2024
consists of 30 competition-grade problems from the 2024 AIME I and II exams. The dataset is formatted for open-ended reasoning with final boxed answers and is widely used in zero-shot evaluations. It focuses on symbolic, algebraic, and geometric reasoning under limited supervision.
AIME 2025
contains 30 problems from AIME I and II 2025, split into Part I and Part II. Compared to 2024, this version includes visual figures (e.g., TikZ-rendered diagrams), making it a more challenging benchmark for symbolic and spatial reasoning tasks. It is commonly evaluated using AveragePass@1.
DeepScaleR-Preview
is a curated subset of a larger multi-agent reasoning benchmark. The dataset emphasizes hierarchical reasoning across multiple steps, with annotations for subgoal structures. It spans mathematical, logical, and planning domains, and is designed to test both compositionality and scalability of reasoning models.
@@
AIME 2024 30 Competition math Olympiad MIT Link
AIME 2025 30 Competition math Olympiad CC BY 4.0 Link
DeepScaleR-Preview [34] Hierarchical reasoning Mixed Apache 2.0 Link
@@ C.5 Reward Design
  • Performing symbolic equivalence checks (e.g., $2x+4$ vs. $4+2x$);
@@ C.7 Environment Setup
For rollout generation, we employ the vLLM backend, which enables efficient and scalable batched decoding with optimized GPU memory usage. Each prompt is decoded to produce five candidate responses, allowing for diverse sampling during training. To ensure stable runtime behavior, we cap GPU memory utilization at 40%. The decoding configuration adopts a stochastic sampling strategy with a temperature of 1.0, top-k=1k=-1 (disabled), and top-$p=1.0$, corresponding to unconstrained nucleus sampling. These settings encourage diverse yet coherent response generation from the model. We constrain the maximum sequence length—comprising both the input prompt and the generated output—to 2048 tokens. Prompts that exceed this limit are filtered out prior to generation, and any attempt to exceed the limit during decoding raises a truncation error. This strict enforcement helps maintain consistency and prevents the introduction of ambiguous or malformed training signals.
@@ Appendix D Lessons Learned from Failed Attempts and Design Iterations
D.1 Offline Data as GRPO-Compatible Rollouts: A Negative Result
@@
D.1.1 Variant I: Direct Injection of Expert Rollouts
@@
D.1.2 Variant II: Self-Rewritten Expert Rollouts
@@
D.2 Conclusion: Offline Rollouts Undermine GRPO Stability
@@
D.3 Offline Data as Few-shots